OpenStudy (anonymous):

What is the inverse of the given function? y = 3x + 9

OpenStudy (anonymous):

I subtract 9 from each side and get y-9=3x divide by 3 on each side, y-9/3=x now what?

OpenStudy (anonymous):

Or am i just totally wrong?

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

Now y and x change roles. Put y for x and x for y.

OpenStudy (anonymous):

I mean the inverse becomes y= (x-9)/3

OpenStudy (anonymous):

inverse of y= (x-9)/3

OpenStudy (anonymous):

I'll give ya another example. Let y = 2x-2 y+2=2x then (y+2)/2 = x This means Inverse of y = (x+2)/2

OpenStudy (anonymous):

Another one . y= (x+2)/2 2y=x+2 then 2y-2 =x This means Invserse of y = 2x-2

OpenStudy (anonymous):

but y=(x-9)/3 isnt any of my answer choices

OpenStudy (anonymous):

A. y= 3x-3 B. y = 1/3x + 3 C. y=3x+3

OpenStudy (agent0smith):

Make sure to use parentheses. y-9/3=x and (y-9)/3 = x are not the same thing.

OpenStudy (anonymous):

y=(x-9)/3 = x/3 - 9/3 = x/3 -3 (B)

OpenStudy (anonymous):

\[y=\frac{ x-9 }{ 3 } = \frac{ x }{ 3 } - \frac{ 9 }{ 3 } = \frac{ x }{ 3} -3 \]

OpenStudy (anonymous):

lol haha cx stupid parentheses cx okay ,,how is it b?

OpenStudy (anonymous):

\[y=\frac{ x }{ 3 } -3 = \frac{ 1 }{ 3 } x - 3 \]

OpenStudy (anonymous):

be had is y= 1/3x+3 not -3

OpenStudy (anonymous):

has *

OpenStudy (anonymous):

Oh! Is there a D ?

OpenStudy (anonymous):

Yeah but its not right

OpenStudy (anonymous):

What is not right?

OpenStudy (anonymous):

D is not right its wrong

OpenStudy (anonymous):

Im guessing they just messed up with the answer choices :P so imma go with b :) thanks

OpenStudy (anonymous):

I beleive so. What we find seems quite right :)